Hi.
I am trying to solve this problem but am confused, because doesn't the 5% change each year as his pay increases each year? Can you help me with a formula?
Joe's career started at age 20, earning $10,000 per year. He received a 5% raise each year and now is retiring at age 52. Show how much Joe will receive each month is his retirement program pays him 75% of the final monthly salary.
Thanks for your help! And Happy Thanksgiving!

Answer by checkley71(8403)

THE FORMULA IS PRINCIPAL(1+RATE)^(NUMBER OF YEARS)
P(1+.05)^32 THE 32 IS THE NUMBER OF YEARS (52-20=32)
10,000(1+.05)^32=$47,649.41 WOULD BE HIS SALARY AFTER 32 YEARS.
.75*47,649.41=$35,737.06 WOULD BE HIS RETIREMENT PAYMENT.